OpenStudy (anonymous):

Solve for x. 4x – (6 – x) = 12 + 4x A. –18 B. 2 C. –2 D. 18

OpenStudy (igreen):

Okay, first distribute -1 into the parenthesis.

OpenStudy (igreen):

-(6 - x) = -1 * 6 + -1 * -x

OpenStudy (igreen):

Can you multiply those? -1 * 6 = ? -1 * -x = ?